OpenStudy (anonymous):

IF 3^X=1/9, WHAT IS THE VALUE OF X?

OpenStudy (anonymous):

hey, you didn't respond to the other problem, are you here?

OpenStudy (anonymous):

\[3^x=\frac{1}{3^2}\rightarrow 3^x=3^{-2}\rightarrow x=-2\]

OpenStudy (anonymous):

YES